### v1

# Fully connected twin-field quantum key distribution network

Abstract
Quantum key distribution (QKD) is a key application in quantum communication, enabling secure key exchange between parties using quantum states. Twin-field (TF) QKD offers a promising solution that surpasses the repeaterless limits, and its measurement-device-independent nature makes it suitable for star-type network architectures. In this work, we propose and demonstrate a fully connected TF-QKD network architecture, where users prepare quantum signals and send them to network nodes. These nodes use splitters to route the signals to measurement units, enabling secure key distribution among arbitrary users. A proof-of-principle demonstration with three users successfully achieved secure key sharing over simulated link losses of up to 30 dB, with an average rate of $5.01\times 10^{-7}$ bit per pulse. Additionally, simulations show that the architecture can support up to $32$ users with a secure key rate of 280.90 bits/s under typical fiber loss conditions. This approach represents a significant advancement in the topology of untrusted-node QKD networks and holds promise for practical, large-scale applications in secure communication.
I Introduction
Quantum key distribution (QKD) is one of the most mature applications of quantum information processing [1]. It uses quantum states to securely distribute symmetric keys between communication parties. Since the first QKD protocol was proposed [2], significant efforts have steadily advanced QKD technology [3, 4, 5, 6, 7, 8]. Recently, a record-breaking achievement was made, with secure key rates exceeding tens of megahertz [5, 6]. Trusted-node QKD networks [9, 10, 11, 12, 13, 14], including integrated space-ground quantum communication networks [14], have been rapidly advancing. However, challenges persist, particularly in long-distance transmission and network expansion, both of which are crucial for the wider deployment of QKD.
Figure 1: (a) Schematic of the physical layer of the fully connected TF QKD network architecture for $N$ users. (b) Schematic of the user transmitter device, which includes frequency-stabilized coherent light based on a local frequency references, intensity modulations, and phase modulations. (c) Schematic of the network node configuration, which includes the same number of polarization alignment modules and $1\times(N-1)$ polarization-maintaining splitters, along with an optical switch array and measurement units satisfying a fully connected setup.
In previous protocols and systems [15, 16], the key generation rate decreases exponentially as the transmission distance increases. Without quantum repeaters, QKD is unable to overcome the fundamental rate-distance limits [17, 18]. Fortunately, twin-field (TF) QKD [19] provides a practical long-distance solution that exceeds the repeaterless limit. Several protocol variants have been proposed to improve practicality [20, 21, 22, 23], such as sending-or-not-sending (SNS) [21] and no-phase-post selection (NPP) [22, 23], which partially eliminate the need for phase-slice post-selection. Significant experimental advances have confirmed TF-QKD’s superior rate-loss scaling [24, 25, 26, 27, 28], with recent demonstrations extending beyond 1000 km [28].
Similar to measurement-device-independent QKD [29], participants in TF-QKD send quantum signals to an intermediate measurement node, enabling immunity to all potential side-channel attacks on measurement devices. This configuration is well-suited for star-type network expansion. However, the stringent requirements of twin-field phase tracking, typically involving additional service fibers and optical frequency-locking hardware, present challenges to building the scalability of TF-QKD networks. Recently, TF-QKD networks based on various TF-QKD network architechture including a simple ring architechture [30] and a $2\times N$ plug-and-play architechture  [31] have been proposed to bypass using phase and frequency locking techonolgy. However, these network architectures require additional components to prevent security vulnerabilities associated with bidirectional optical paths [32, 33, 34, 35]. Furthermore, to enable multi-user key distribution, time-division multiplexing of a single detection unit is required. This presents a significant challenge to the detector’s bandwidth as the number of users increases, thereby reducing the achievable secure key rate per user.
In this study, we propose and demonstrate a fully connected TF-QKD network architecture. Our architecture is inspired by recent advancements in protocols [36, 37, 38] and open-architecture TF-QKD schemes [39, 40, 41, 42, 43], eliminating the need for additional channels and devices to synchronize the frequencies of independent light sources, thereby offering greater flexibility and scalability. Furthermore, each user prepares quantum signals and sends them to a network node in single-path transmission, which inherently provides immunity to bidirectional transmission security vulnerabilities. Furthermore, the node employs beam splitters to route the received quantum signals to different single-photon interference measurement units, enabling secure key distribution among arbitrary users. We conduct a proof-of-principle demonstration of a three-user fully connected TF-QKD network, where secure key sharing is successfully achieved between any pair of users over a simulated link loss of up to $30$ dB, with an average rate of $5.01\times 10^{-7}$ bit per pulse. Additionally, we perform simulations based on the parameters of the experimental system. The results show that our network architecture can support up to $32$ users in a symmetric scenario, with a secure key rate of $280.90$ bits/s between user pairs under a fiber channel loss of $20$ dB. This network architecture enhances the topology of untrusted-node QKD networks and promotes their practical applications.
II Network architecture
The fully connected TF QKD network architecture is depicted in Figure 1. The physical layer adopts a star topology consisting of end users, fiber channels, and untrusted network node, as shown in Figure 1(a). Each user is equipped with a transmitter, located at the terminal of channel extending from the central node, as shown in Figure 1(b). The network node includes polarization monitors, polarization-maintaining splitters, an optical switch array, and single-photon interference measurement units. Full connectivity among multiple users is achieved through the designed of nodal components, as exemplified in Figure 1(c).
Figure 2: Schematic of our experimental setup. A commercial laser serves as the light source for all users. Each frame of light pulse is modulated by the first intensity modulation module (IM-$1$) into $400$ signal pulses, $600$ reference pulses, and $24$ vacuum pulses. The intensity modulation module includes beam splitter (BS) and a phase modulator (PM). Subsequently, each user employs a second intensity modulation module (IM-$2$) and a phase modulator (PM) to implement encoding, decoy intensity modulation, and discrete random phase modulation on the optical pulses. The pulses are then attenuated via a variable optical attenuator (VOA), followed by a simulated channel using VOA. At the network node, quantum signals from user pass through polarization controller (PC) and polarization beam splitter (PBS), then couple into a polarization-maintaining splitter (PMS). The signals from the user pairs interfere at a BS within the measurement unit (MU). Finally, the quantum signals are detected by superconducting nanowire single-photon detectors (SNSPDs).
An $N$-user network instance is considered here to further elucidate our network. Each user’s transmitter comprises a laser source, intensity modulators, and phase modulators. The user prepares frequency-stabilized light signal based on a local frequency reference, then modulate the light signal’s intensity and phase according to the specific TF-QKD protocol. The modulated signal is transmitted to network nodes via fiber channel.
The polarization monitor at the network node tracks and compensates for polarization drift in the user signal. The signal is then aligned and coupled into a $1\times(N-1)$ polarization-maintaining splitter, where it is randomly directed to one of its branches. The $N-1$ branches of the splitter enable connections between the user and the remaining $N-1$ users. The $(N^{2}-N)\times(N^{2}-N)$ optical switch array allocates input ports to the branches of all splitters. For example, any user who intends to communicate with User-11{1}1 can establish a complete quantum channel by accessing the measurement unit through the optical switch. For an $N$-user network, up to $N\times(N-1)/2$ measurement units are required to achieve full connectivity. Figure 1(c) shows an $N$-user fully connected configuration. Considering some dedicated line communication scenarios, network node can also form a static configuration by directly wiring splitters and measurement units, as demonstrated in the three-user network in Section III.
In this network, key distribution between any pair of users is implemented based on the TF-QKD protocol, which requires coherent control of the light fields from remote users. The frequency difference between independent user lasers or phase shifts induced by channel fluctuations affect the coherence of these light fields. Building on recent advancements [43], users can employ acetylene-stabilized laser, eliminating the need for a shared optical frequency reference. This makes it possible for the independent lasers of users in the network to maintain frequency alignment using a local frequency reference. Furthermore, there are mature solutions to effectively compensate for the phase shifts caused by fiber fluctuations over several hundred kilometers, either through real-time feedback [26, 27, 39] or post-processing techniques [28, 43].
III Experiment
To validate the proposed architecture, we build a proof-of-principle experimental setup, which includes three fully connected users in a TF-QKD network. Figure 2 shows the schematic of experimental setup. A laser source operating at $1550$ nm generates optical pulses at a $100$ MHz clock rate, with a pulse width of approximately $500$ ps. These pulses enter the first intensity modulation module (IM-$1$), which is a Sagnac-based interferometer. The optical pulses are modulated into sequential frame-based signals. Specifically, each frame consists of $1024$ optical pulses, comprising $400$ signal pulses, $600$ reference pulses and $24$ vacuum pulses. These pulses are then sent to the three users.
Each user encodes the signal pulses according to the three-intensity sending-or-not-sending (SNS) TF QKD protocol [44, 45] by employing the second intensity modulation module (IM-$2$), thereby producing a signal window for key extraction and a decoy window to estimate information leakage. Subsequently, a phase modulator (PM) is employed to perform $16$-level random phase modulation over a range of $2\pi$ to the signal pulses. Finally, the signal pulses are attenuated to the single-photon level using a variable optical attenuator (VOA) before entering the channel.
Figure 3: Interference results between users in a three-user fully connected network within $200$ ms. The interference results are captured by recording the counts of the detectors D0 and D1. The average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively.
Signals from different users reach the network node through a channel simulated by a VOA. The arriving quantum signals are first processed by a polarization controller (PC) to align their polarization to the one of orthogonal modes of the followed polarization beam splitter (PBS). The signals are then efficiently coupled into a $1\times 2$ fast-axis block polarization-maintaining splitter (PMS), and randomly routed to different branches with the same polarization. The signals from the user pairs interfere with the same polarization at the beam splitter (BS) within the measurement unit (MU). The interference outputs between user pairs are detected by two superconducting nanowire single-photon detectors (SNSPDs), with detection efficiency around $45\%$ and a dark count rate of about $8\times 10^{-8}$. The detected photon events are recorded by the TDC and then processed by a computer. The detection events of the reference pulse are used in post-processing to estimate the phase fluctuation between users. Within the effective arrival time $t$, the detection events in the decoy window are further filtered according to the following condition:
$\left|\cos\left(\theta_{i}-\theta_{j}-\varphi_{ij}\right)\right|\geq\cos(\pi/16),$ (1)
where $\theta_{i}$ ($\theta_{i}$) is the encoded phase of user-$i$ ($j$) to be announced in post-processing, and $\varphi_{ij}$ is the phase fluctuation between users. We also apply the actively odd-parity pairing (AOPP) method [45] to suppress bit-flip errors and improve the secure key rate.
IV Results
Using the described network system, we first measure the interference between users to preliminarily assess system performance. The measurement is conducted with the phase modulator installed but not actively modulating. The interference results are obtained by monitoring the counts from detectors in the MU, as shown in Figure 3. We observe that the average interference visibilities for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ are $90.87\%$, $95.82\%$, and $96.72\%$, respectively, over a period of approximately $200$ ms. Due to the single laser scheme adopted, phase fluctuations affecting interference visibility in the experiment primarily originate from optical components.
We then perform a three-user TF-QKD network experiment. The transmittance for the users is simulated using optical attenuators, which model the transmittance of fiber links with lengths of $100$ km and $150$ km, assuming a loss coefficient of $0.2$ dB/km. Based on the similar transmittance of all users, we select the same implementation parameters for simplicity. Considering the finite-key effect, the secure key rate of after AOPP is calculated using  [45]
$R=$ $\frac{1}{N}\{{n_{1}^{\prime}(1-h(e_{1}^{\prime ph}))-fn_{t}^{\prime}h(E^{\prime})}$ (2)
${-2\log_{2}\frac{2}{\varepsilon_{cor}}-4\log_{2}\frac{1}{\sqrt{2}\varepsilon_{PA}\hat{\varepsilon}}}\},$
where $N$ is the total number of send pulse pairs, $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ is the Shannon entropy function. $f$ is the error correction efficiency equal to $1.1$. $n_{t}^{\prime}$ is the number of remaining bits after AOPP. $E^{\prime}$ is the bit-flip error rate of the remaining bits after AOPP. $\varepsilon_{cor}$ is the failure probability of error correction, $\varepsilon_{PA}$ is the failure probability of privacy amplification, $\hat{\varepsilon}$ is the coefficient obtained using the chain rules of smooth min- and max-entropy, with all three parameters being set to $10^{-10}$. Detailed parameter estimation for the three-intensity SNS-TF QKD with AOPP protocol can be found in Appendix A.
Figure 4: (a) Experimental results in a three-user fully connected network. (b) The secure key rate per user pair in a $32$-user fully connected network.
For each transmittance level, a total of $10^{10}$ signal pulses are emitted by each user. The detection events of the reference pulses are extracted within a time window of $200$ ms, and only the data with stable interference are retained. We estimate phase fluctuation from the reference data during this time interval and apply these values to compensate for the phase fluctuation of signal pulses for secure key generation. As shown in Figure 4(a), the secure key rate for user pairs $1$-$2$, $1$-$3$, and $2$-$3$ is respectively $2.38\times 10^{-8}$ ($6.87\times 10^{-6}$), $1.88\times 10^{-7}$ ($9.35\times 10^{-6}$), and $1.29\times 10^{-6}$ ($2.02\times 10^{-5}$) bits/pulse under a simulated channel loss of $30$ ($20$) dB. Detailed implementation parameters and measurement results are summarized in Appendix B.
To evaluate the network’s performance with an increasing number of users, we consider our system parameters along with the splitter loss and clock rate parameters from reference [46]. The network performance in a symmetric scenario is shown in Figure 4(b). The results indicate that this network architecture can support a $32$-user fully connected network, achieving secure key rates of $490.63$ bits/s per user pair for a simulated fiber loss of $10$ dB and $208.90$ bits/s per user pair for a simulated fiber loss of $20$ dB.
V Conclusion
In summary, we propose a fully connected TF-QKD network architecture, designed to efficiently support secure key distribution among multiple users without relying on trusted nodes. Each user is connected to a central node equipped with splitters, which randomly route quantum signals to measurement units, enabling secure key sessions between arbitrary user pairs. A three-user experimental demonstrate the feasibility of this network expansion strategy, achieving stable secure key distribution across all pairwise connections. Furthermore, simulations show that this network can support up to $32$ users with a secure key rate of $280.90$ bits/s for user pairs under a simulated channel loss of $20$ dB.
This work advances the practical deployment of untrusted-node QKD networks, offering a scalable and efficient solution for future quantum-secured communication infrastructures. Furthermore, this architecture, combined with compact integrated chip-based systems [47] and cost-effective hybrid quantum-classical system [48], is expected to further drive the practical application of QKD network. Beyond QKD, this architecture could be adapted into other quantum protocols, such as quantum digital signature [49] and quantum blockchain [50].
Acknowledgment
This study was supported by the National Natural Science Foundation of China (Nos. 62171144 and 62031024), Guangxi Science Foundation (Nos. 2025GXNSFAA069137 and GXR-1BGQ2424005), and Innovation Project of Guangxi Graduate Education (No. YCSW2022040).
Appendix A Three-intensity SNS-TF QKD with active odd parity pairing (AOPP) protocol
Table 1: The three-intensity parameters under different simulated losses, including intensities and probabilities. $\{\mu_{o},\mu_{x},\mu_{y}\}$ represent three sources with different intensities. $p_{x}$ is the probability of sending decoy intensity $\mu_{x}$. $p_{y}$ is the probability of sending the signal window. $\varepsilon$ is the probability of sending during the signal window. $\mu_{ref}$ is the intensity of reference pulse.
$Loss$ (dB) $\mu_{y}$ $\mu_{x}$ $\mu_{o}$ $p_{x}$ $p_{y}$ $\varepsilon$ $\mu_{ref}$
$20$ $0.44$ $0.01$ $0.0016$ $0.23$ $0.72$ $0.25$ $1.5$
$30$ $0.43$ $0.01$ $0.0016$ $0.36$ $0.53$ $0.25$ $1.5$
Here, we describe the three-intensity SNS-TF QKD protocol [21, 44] adopted in our network demonstration, where the AOPP method is employed to extract the final secret key. Given that user-$i$ and user-$j$ in the network request key distribution session. User-$i(j)$ possesses three sources with different intensities, denoted as $\mu_{o_{i}},\mu_{x_{i}},\mu_{y_{i}}(\mu_{o_{j}},\mu_{x_{j}},\mu_{y_{j}})$, where $\mu_{o_{i}}(\mu_{o_{j}})$ represents vacuum source. In each time window, user-$i(j)$ randomly determines whether the optical pulse originates from the vacuum source $\mu_{o_{i}}(\mu_{o_{j}})$ with probability $p_{o}$, the decoy source $\mu_{x_{i}}(\mu_{x_{j}})$ with probability $p_{x}$, or a signal window with probability $p_{y}=1-p_{o}-p_{x}$. If a signal window is determined, user-$i(j)$ then randomly chooses between the source $\mu_{o_{i}}(\mu_{o_{j}})$ with probability $1-\epsilon$ and the source $\mu_{y_{i}}(\mu_{y_{j}})$ with probability $\epsilon$. For vacuum source in the signal window, user-$i(j)$ encodes bit $0(1)$, whereas for source $\mu_{y_{i}}(\mu_{y_{j}})$ in the signal window, user-$i(j)$ encodes bit $1(0)$.
Considering a key distribution session, the total number of pulses sent by user-$i(j)$ is $N$. The number of source pairings $lr(l\in\{\mu_{o_{i}},\mu_{x_{i}},\mu_{y_{i}}\},r\in\{\mu_{o_{j}},\mu_{x_{j}},\mu_{y_{j}}\})$ can be represented as $N_{lr}$, and the total number of corresponding single detector response events is represented as $n_{lr}$. The yield of source $lr$ can be defined as $S_{lr}=n_{lr}/N_{lr}$, and the expected value is represented as $\left\langle S_{lr}\right\rangle$. Based on the above definitions, we have:
$\begin{array}[]{c}N_{\mu_{o_{i}}\mu_{o_{j}}}=\left[p_{o}^{2}+2p_{o}p_{y}(1-\varepsilon)\right]N,\\ N_{\mu_{o_{i}}\mu_{x_{j}}}=N_{\mu_{x_{i}}\mu_{o_{j}}}=\left[p_{o}+p_{y}(1-\varepsilon)\right]p_{x}N,\\ N_{\mu_{o_{i}}\mu_{y_{j}}}=N_{\mu_{y_{i}}}\mu_{o_{j}}=p_{o}p_{y}\varepsilon N.\end{array}$ (3)
For the time window in which user-$i$ and user-$j$ select source $\mu_{x_{i}}$ and $\mu_{x_{j}}$, the users will announce the phases encoded and perform post-selection on the single detector response events based on the following conditions:
$1-\left|\cos\left(\theta_{i}-\theta_{j}-\varphi_{ij}\right)\right|\leq\lambda,$ (4)
where $\theta_{i}$ and $\theta_{j}$ are the encoding phase of the pulses for user-$i$ and user-$j$ respectively, $\varphi_{ij}$ the phase fluctuation between user-$i$ and user-$j$, and $\lambda$ is a small positive value. Let the number of source pairings satisfying the above condition be denoted as $N_{x}$, and the number of error events as $m_{x}$. Then, the error rate for the window $\mu_{x_{i}}\mu_{x_{j}}$ us given by $T_{x}=m_{x}/N_{x}$.
To obtain the secure key rate, the users need to estimate the lower bound of the number of untagged bits and the upper bound of the phase error rate of the untagged bits. After performing decoy state analysis, the lower bound of the expected counting rate for the states $|01\rangle$$\langle 01|$ and $|10\rangle$$\langle 10|$ can be expressed as follows:
$\underline{\left\langle s_{01}\right\rangle}$ $=\frac{\mu_{y}^{2}e^{\mu_{x}}\underline{\left\langle S_{\mu_{o_{i}}\mu_{x_{j}}}\right\rangle}-\mu_{x}^{2}e^{\mu_{y}}\overline{\left\langle S_{\mu_{o_{i}}\mu_{y_{j}}}\right\rangle}-\left(\mu_{y}^{2}-\mu_{x}^{2}\right)\overline{\left\langle S_{\mu_{o_{i}}\mu_{o_{j}}}\right\rangle}}{\mu_{y}\mu_{x}\left(\mu_{y}-\mu_{x}\right)},$ (5)
$\underline{\left\langle s_{10}\right\rangle}$ $=\frac{\mu_{y}^{2}e^{\mu_{x}}\underline{\left\langle S_{\mu_{x_{i}}\mu_{o_{j}}}\right\rangle}-\mu_{x}^{2}e^{\mu_{y}}\overline{\left\langle S_{\mu_{y_{i}}\mu_{o_{j}}}\right\rangle}-\left(\mu_{y}^{2}-\mu_{x}^{2}\right)\overline{\left\langle S_{\mu_{o_{i}}\mu_{o_{j}}}\right\rangle}}{\mu_{y}\mu_{x}\left(\mu_{y}-\mu_{x}\right)},$
then the lower bound of the counting rate for the untagged bits can be expressed as follows:
$\underline{\left\langle s_{1}\right\rangle}=\frac{1}{2}(\underline{\left\langle s_{01}\right\rangle}+\underline{\left\langle s_{10}\right\rangle}).$ (6)
The lower bound of the expected values for the number of untagged bit $1$ and untagged bit 00 are given by:
$\underline{\left\langle n_{10}\right\rangle}=Np_{y}^{2}\epsilon(1-\epsilon)\mu_{y}e^{-\mu_{y}}\underline{\left\langle s_{10}\right\rangle},$ (7)
$\underline{\left\langle n_{01}\right\rangle}=Np_{y}^{2}\epsilon(1-\epsilon)\mu_{y}e^{-\mu_{y}}\underline{\left\langle s_{01}\right\rangle}.$
The upper bound of the expected values for the phase error rate $1$ is given by:
$\overline{\left\langle e_{1}^{ph}\right\rangle}=\frac{\left\langle T_{x}\right\rangle-e^{-2\mu_{x}}\left\langle S_{\mu_{o_{i}}\mu_{o_{j}}}\right\rangle/2}{2\mu_{x}e^{-2\mu_{x}}\underline{\left\langle s_{1}\right\rangle}}.$ (8)
Table 2: Detailed experimental results under different simulated losses, including the secure key rate considering finite-key effect and AOPP. The detection events are labeled as ”Detected-$ijab$”, where ’$i(j)$’ and ’$a(b)$’ represent the signal $Z$ or decoy $X$ time windows and intensities determined by user-$i(j)$.
User1111-$2$ User1111-$3$ User2222-$3$ User1111-$2$ User1111-$3$ User2222-$3$
Loss (dB) 20 30
$N$ $10^{10}$ $10^{10}$
$R$ (bits/pulse) $6.8\times 10^{-6}$ $9.35\times 10^{-6}$ $2.02\times 10^{-5}$ $2.38\times 10{-8}$ $1.88\times 10^{-7}$ $1.29\times 10^{-6}$
Detected $ZZyy$ 3984035 4077966 3645054 854517 601800 648818
Detected $ZZoy$ 6088650 5973265 5217181 1022963 954205 912414
Detected $ZZyo$ 5160981 4848325 6250619 912024 751293 1024881
Detected $ZZoo$ 147816 85275 105192 12931 11578 16709
Detected $ZXyx$ 2348664 2261991 2816486 791683 1046887 1003973
Detected $ZXox$ 166084 182167 213010 95600 32038 71703
Detected $ZXyo$ 445273 594330 552265 179775 359825 206623
Detected $ZXoo$ 14484 6656 12434 5705 3344 6423
Detected $XZxy$ 2421535 1853833 2490897 992859 983220 784572
Detected $XZoy$ 587264 501135 407128 349189 202943 228031
Detected $XZxo$ 211380 192685 140001 34267 97368 70640
Detected $XZoo$ 15243 7888 12356 7475 2976 4317
Detected $XXxx$ 130416 180341 127632 103282 104274 145071
Detected $XXox$ 16158 13966 24151 23824 12949 22445
Detected $XXxo$ 18007 14983 12346 13302 25180 21872
Detected $XXoo$ 350 788 475 702 1518 651
Detected $XXxx$-accepted 19109 26176 20232 17799 19131 23062
Correct $XXxx$-accepted 17381 24176 18694 16538 17859 21208
QBER $ZZ$ before AOPP 26.86% 27.78% 24.64% 30.95% 26.45% 25.57%
QBER $ZZ$ after AOPP 1.84% 1.19% 1.16% 1.16% 0.95% 1.13%
We further calculate the lower bounds on the number of untagged bit $n_{1}^{\prime}$ and the phase error rate $e_{1}^{\prime ph}$ after AOPP according to the methods in reference [45]. The related formulas of $n_{1}^{\prime}$ are as follows:
$u$ $=\frac{n_{g}}{2n_{odd}},$ (9)
$\underline{n_{10}}$ $=\varphi^{L}(\underline{\langle n_{10}\rangle}),$
$\underline{n_{01}}$ $=\varphi^{L}(\underline{\langle n_{01}\rangle}),$
$\underline{n_{1}}$ $=\underline{n_{01}}+\underline{n_{01}},$
$n_{1}^{r}$ $=\varphi^{L}\left(\frac{\underline{n_{1}}}{n_{t}}\frac{\underline{n_{1}}}{n_{t}}\frac{un_{t}}{2}\right),$
$n_{10}^{\prime}$ $=2n_{1}^{r}\left(\frac{\underline{n_{10}}}{\underline{n_{1}}}-\sqrt{\frac{-\ln\epsilon}{2n_{1}^{r}}}\right),$
$n_{01}^{\prime}$ $=2n_{1}^{r}\left(\frac{\underline{n_{01}}}{\underline{n_{1}}}-\sqrt{\frac{-\ln\epsilon}{2n_{1}^{r}}}\right),$
$n_{min}$ $=\min(n_{01}^{\prime},n_{10}^{\prime}),$
$n_{1}^{\prime}$ $=2\varphi^{L}\left(n_{min}\left(1-\frac{n_{min}}{2n_{1}^{r}}\right)\right),$
where $n_{t}$ is the raw key generated by users-$i$ and user-$j$, $n_{g}$ is the number of pair if users-$i$ and user-$j$ perform AOPP to their raw keys. $n_{odd}$ is the number of pair with odd-parity when user-$j$ randomly groups their raw keys two by two. $\epsilon$ is the failure probability of parameter estimation. $\varphi^{U}(x)$ and $\varphi^{L}(x)$ are the upper and lower bounds after using Chernoff bound [51] to estimate the real values based on the expected values.
The related formulas of $e_{1}^{\prime ph}$ are as follows:
$r$ $=\frac{\underline{n_{1}}}{\underline{n_{1}}-2n_{1}^{r}}\ln\frac{3(\underline{n_{1}}-2n_{1}^{r})^{2}}{\epsilon},$ (10)
$e_{\tau}$ $=\frac{\varphi^{U}(2n_{1}^{r}\overline{\langle e_{1}^{ph}\rangle})}{2n_{1}^{r}-r},$
$\overline{M_{s}}$ $=\varphi^{U}((n_{1}^{r}-r)e_{\tau}(1-e_{\tau}))+r,$
$e_{1}^{\prime ph}$ $=\frac{2\overline{M_{s}}}{n_{1}^{\prime}}.$
Finally, the secure key rate can expressed by:
$R=$ $\frac{1}{N}\{{n_{1}^{\prime}(1-h(e_{1}^{\prime ph}))-fn_{t}^{\prime}h(E^{\prime})}$ (11)
${-2\log_{2}\frac{2}{\varepsilon_{cor}}-4\log_{2}\frac{1}{\sqrt{2}\varepsilon_{PA}\hat{\varepsilon}}}\}.$
Appendix B Detailed experimental results
Here, we summarize the parameters of the three-intensity SNS protocol used in our experiment, as shown in Table 1. Table 2 presents detailed experimental results under different simulated losses, including the single-detector response events used to calculate the secure key rate and the AOPP results.